# FRABench and GenEval: Scaling Fine-Grained Aspect Evaluation across Tasks, Modalities

### v2

@@ Appendix A Aspect Sources and Definitions
NLG (Public) Summarization Train 218 IU (Public) Detailed Image Captioning Train 2.1k
@@
Translation Test 130 ITIG (Generated) Storytelling Generation Test 50
@@
Table 9: The detailed statistical information of the sub-tasks is not shown in Figure 3
@@
IG Text-to-Image Generation Show-o-1.3B, Seed-X-17B, Flux-12B, Stable-Diffusion-3.5-Large, Ground Truth
@@
ITIG Visual Story Completion miniGPT, GILL, MM-Interleaved, GPT4o+SDXL, Ground Truth
@@
Table 10: Different LMMs used for generating pairwise responses for sub-tasks under the “Generated” column in Figure 3
Appendix B Analysis of the FRABench
In this section, we provide detailed information and comprehensive analysis about the FRABench. The dataset is currently available via an anonymous link: https://github.com/GenEval-1/GenEval.
B.1 More Detailed Information about the Dataset
We provide the detailed information about sub-tasks for NLG and IU in Table 9, as this information is not shown in Figure 3. For the tasks in the “Generated” column in Figure 3, due to the absence of pairwise data, we use four LMMs with varying performance levels to generate pairwise responses for each query, as illustrated in Table 10.
B.2 Aspect Diversity Analysis
Following prior work [31, 44], we analyze the ROUGE-L distribution between aspects in the training set and the FRA-OOD, respectively. Specifically, we sample pairs of aspects from the training set and compute their ROUGE-L scores. The overall distribution is shown in Figure 7, which indicates that the selected aspects are distinct from one another, confirming the inclusion of diverse and novel aspects in the training set. Additionally, we measure the ROUGE-L scores between aspects by sampling one aspect from the training set and another from the FRA-OOD . As shown in Figure 8, aspects in the training set and those in the FRA-OOD do not overlap, demonstrating the validity of our out-of-domain evaluations.
B.3 Position Bias Analysis
We provide the counts of evaluation labels in the training set, FRA-ID, FRAUAs-OOD, and FRA-OOD, showing cases where Response 1 is better than, equal to, or worse than Response 2, as illustrated in Figures 10 to 12. As shown in Figure 10, the four tasks in the training set maintain a relatively consistent distribution. In Table 11, we also evaluate the position consistency of GenEval on the FRA-ID and FRA-OOD. We perform two inferences on the same sample, swapping the positions of the responses in the second inference, and calculate the average consistency, i.e., whether the results from the two inferences align. These experimental results demonstrate that GenEval achieves higher positional consistency compared to other baselines.
@@ B.3 Position Bias Analysis
Figure 9: Statistics of pairwise evaluation labels in the training set.
Figure 10: Statistics of pairwise evaluation labels in the FRAUAs-OOD.
Figure 11: Statistics of pairwise evaluation labels in the FRA-ID.
Figure 12: Statistics of pairwise evaluation labels in the FRA-OOD.
B.4 Aspect Distribution Across Different Sub-tasks
We provide detailed information on the aspects selected for the training set, FRA-ID, FRAUAs-OOD, and FRA-OOD , as presented in Table 12. Additionally, we include statistics regarding the number of evaluation labels used for training and testing in each sub-task. Blue-marked aspects are human-annotated, while the rest are annotated by GPT-4o. The human-annotated data in IG is sourced from the ImageRewardDB dataset [10].
@@ B.4 Aspect Distribution Across Different Sub-tasks
Training Set & FRA-ID Summarization Train / Test Clarity, Grammaticality, Coherence, Fluency Coverage, Length Constraint, Layout, Conciseness 1.7k / 0.4k
@@
General Communication Train / Test Gender Bias, Ethnic Bias, Regional Bias Accuracy, Information Richness 11.8k / 2.6k
@@
General Visual Conv Train / Test Clarity, Harmfulness Helpfulness, Instruction Following 8.4k / 1.6k
Text-rich Und Train / Test Harmfulness, Vocabulary Complexity, Clarity Accuracy, Instruction Following 10.5k / 2k
Robustness-oriented Instructions Train / Test Spelling Accuracy, Structure Accuracy, Privacy Violations Instruction Inconsistency, Context Inconsistency 19.6k / 2.5k
Medical Image Und Train / Test Vocabulary Complexity, Bias, Sentence Complexity Accuracy 8.3k / 2k
@@
Simple Image Captioning Train / Test Readability, Bias Conciseness, Object Alignment 12.5k / 2.4k
@@
Detailed Image Captioning Train / Test Coherence, Fluency, Clarity, Bias Completeness, Attribute Alignment, Spatial Alignment 27.2k / 5.2k
@@
Text-to-Image Generation (COCO&DB) Train / Test Harmfulness, Fidelity Count Alignment, Object Relationship 49.2k / 7.5k
@@
Visual Story Completion Train / Test Scene Consistency, Bias, Toxicity, Character Consistency Creativity, Engagingness, Completeness 39.8k / 7.7k
@@
Multimodal Script Generation Train / Test Action Consistency, Coherence, Image Repetition Completeness, Feasibility 40.5k / 6.2k
@@
FRAUAs-OOD Summarization Test Sentence Complexity, Vocabulary Complexity, Readability - 1.3k
@@
General Communication Test Readability, Harmfulness, Sentence Complexity - 5.4k
@@
General Visual Conv Test Fluency, Coherence, Grammaticality, Readability - 1.6k
Text-rich Und Test Fluency, Coherence, Grammaticality, Readability - 1.6k
Robustness-oriented Instructions Test Readability, Bias, Fluency, Coherence - 1.6k
Medical Image Und Test Fluency, Coherence, Readability, Grammaticality - 1.6k
Simple Image Captioning Test Fluency, Clarity, Ethnic Bias, Coherence, - 2.4k
@@
Detailed Image Captioning Test Readability, Grammaticality - 1.6k
@@
Visual Story Completion Test Readability, Coherence, Grammaticality - 3.8k
@@
Multimodal Script Generation Test Grammaticality, Fluency - 2.4k
@@
FRA-OOD Question Generation Test Clarity, Grammaticality, Fluency, Coherence Engagement, Difficulty 1.1k
@@
Academic VQA Test Fluency, Readability, Vocabulary Complexity Expertise, Correctness, Explainability 3.6k
@@
Medical VQA Test Sentence Complexity, Readability, Vocabulary Complexity Terminology, Transparency 3.6k
@@
Embodied Decision-makin Test Coherence, Clarity, Grammaticality, Gender Bias Interpretability, Completion, Riskiness 4.8k
@@
Text-to-Image Generation (GenAI-Bench) Test Harmfulness, Fidelity Count Alignment, Object Relationship 6.7k
@@
Activitynet Continuation Test Readability, Harmfulness, Coherence, Semantic Consistency Reasonableness, Fulfillment 0.5k
@@
Storytelling Generation Test Readability, Harmfulness, Scene Consistency, Image Coherence Originality, Simplification 0.6k
@@
Table 12: Aspects and evaluation label counts in the Training Set, FRA-ID, FRAUAs-OOD, and FRA-OOD. Blue-marked aspects are human-annotated, while the rest are annotated by GPT-4o.
@@
Bias Bias  [45, 46, 47, 46] T, ITI Bias evaluates the presence and extent of gender, regional, and cultural bias in generated content. It examines whether the content reinforces stereotypes, shows favoritism or discrimination toward specific regions, or reflects racial or ethnic prejudice. The assessment focuses on language, tone, and representation, aiming to identify harmful biases or imbalanced portrayals.
@@
Gender Bias GenB  [46, 48, 49, 49] T, ITI This criterion evaluates whether the generated content reinforces traditional gender stereotypes or presents biased views based on gender. It examines whether certain roles, traits, or behaviors are unfairly associated with specific genders, such as linking nurturing to women or leadership to men.
@@
Cultural Bias CulB  [46] T, ITI This criterion evaluates the model’s understanding of and respect for cultural diversity. It assesses whether the generated text appropriately acknowledges cultural differences and avoids reinforcing harmful generalizations or stereotypes about specific cultures. For instance, models may misinterpret culturally specific idioms or expressions, or fail to recognize regional language variants, which can result in outputs that lack cultural sensitivity or misrepresent certain communities.
@@
Regional Bias RegB  [46, 49] T, ITI This criterion evaluates whether the generated content shows a bias toward or against any particular nationality or region. It ensures that all national or regional identities are treated fairly, without favoritism, discrimination, or negative stereotyping.
@@
Ethnic Bias EthB  [50, 48, 49] T, ITI This criterion assesses whether the generated content reflects racial or ethnic biases, either by perpetuating stereotypes or marginalizing certain groups. It looks for whether certain ethnic or racial groups are portrayed negatively or unfairly advantaged, reinforcing harmful biases.
@@
Toxicity Tox  [51, 47, 52] T, ITI This criterion evaluates whether the model avoids generating harmful, violent, explicit, or inappropriate content. It ensures the system does not produce material that could cause emotional harm, violate legal norms, or infringe on privacy. This includes preventing the generation of discriminatory, NSFW, or harmful language and images.
@@
Harmfulness Har  [53, 47, 52] T, ITI This criterion evaluates whether the generated content promotes, glorifies, or depicts any form of violence, abuse, or harm—be it physical, emotional, or psychological. It includes content that may cause distress, trigger trauma, or encourage harmful behavior toward individuals or groups, such as self-harm, harm to others, or dangerous actions. Images that cause psychological discomfort when seen are harmful.
@@
Privacy Violation Priv  [54, 47, 52] T, ITI This criterion evaluates whether the generated content reveals personal, private, or sensitive information about individuals without consent, violating privacy norms or laws. This includes disclosing private details such as names, locations, health conditions, or other confidential information. The content should respect an individual’s privacy rights, in accordance with legal and ethical standards.
@@
Readability Rea  [55, 15, 9, 56] T, ITI Readability measures how easily a text is understood by its intended audience. It focuses on overall fluency and the clarity of ideas to ensure they are presented simply and align with the readers’ comprehension levels. It evaluates the quality of both inter-sentences and intra-sentences, checking if they are grammatically correct, naturally written, with clear meanings, and good context-relatedness and coherence, allowing readers to absorb information effortlessly.
@@
Coherence Coh  [5, 15, 9] T, ITI Coherence refers to the logical consistency and interconnection of ideas within the text. It ensures that the text follows a clear, organized progression, with each section logically linking to the next. A coherent text allows the reader to easily follow the author’s argument or narrative, with smooth transitions and no sudden shifts in topic or gaps in reasoning.
@@
Clarity Cla  [9, 39] T, ITI Clarity measures how easily the text can be understood. It emphasizes the use of straightforward, unambiguous language, avoiding complex vocabulary and convoluted sentence structures that could confuse the reader.
Text Repetition TexR  [9, 57] T, ITI Text repetition refers to the unnecessary duplication of phrases, sentences, or ideas that do not add new meaning. Repetition disrupts the flow, reduces clarity, and can make the text feel redundant.
Sentence Complexity SenC  [58, 59] T, ITI This criterion evaluates how sentence structures affect readability. Complex or overly intricate sentences may hinder comprehension, while simpler structures enhance clarity.
@@
Vocabulary Complexity VocC  [58, 59] T, ITI This criterion evaluates whether the vocabulary used in the text is well-suited to the target audience’s comprehension level. The language should prioritize clarity by avoiding overly complex or obscure terminology that may hinder understanding, while also ensuring it is not excessively simplistic for the intended readers.
@@
Fluency Flu  [9, 39, 5, 15] T, ITI Fluency measures the smoothness of the language and sentence construction, ensuring the text flows seamlessly and is grammatically correct. The fluent text flows naturally, uses familiar phrasing, and avoids awkward or forced transitions.
@@
Grammaticality Gra  [9, 15] T, ITI Grammaticality measures how well the text adheres to the rules of grammar, including sentence structure, verb tense, subject-verb agreement, and punctuation.
@@
Spelling Accuracy SpeA  [9, 15] T, ITI This criterion assesses the correctness of spelling in the text, ensuring there are no typographical errors or misspelled words.
@@
Tense Consistency TenC  [9, 15] T, ITI This criterion assesses whether the sentence maintains consistent tense usage (past, present, future) throughout. The tense should remain uniform unless a change is contextually required, such as when indicating a change in the time frame.
@@
Structure Accuracy StrA  [9, 15] T, ITI This criterion evaluates the grammatical accuracy of sentences, focusing on the proper alignment of subject-verb-object and the correct linkage of main and subordinate clauses. It ensures that sentences are constructed following standard grammatical rules, which enhances clarity and readability.
@@
Image Quality ImaQ  [60, 10, 11] I, ITI Image quality primarily evaluates the fidelity of the generated image, ensuring it accurately represents the described objects and their characteristics. It also assesses whether the image contains any harmful elements, such as offensive content, misleading representations, or harmful stereotypes.
@@
Fidelity Fid  [10, 11] I, ITI Fidelity evaluates how accurately the generated image represents the shape, characteristics, and behavior of the objects described, ensuring they align with real-world expectations. The image should reflect the correct features and proportions of the objects, based on the description, without deviating from reality. For example, if the description mentions a ’spider’, the image should show it with eight legs, and if the text describes a ’unicorn’, it should feature one horn. Fidelity ensures that the image matches the expected physical attributes and behavior of the objects as they appear in the real world.
@@
Aesthetic Aes  [61, 11] I, ITI Aesthetic measures how pleasing and emotionally striking an image is. It checks if the image combines elements like color and light to create a visually appealing experience that resonates with viewers. This measure helps assess whether the image captures beauty and evokes an emotional response, making it memorable and impactful.
@@
Image Coherence ImaC  [60] ITI Image coherence measures the consistency in style and subject representation across images. This includes textures, color palettes, lighting, rendering styles, and maintaining consistent physical attributes, clothing, and behavioral traits. Image coherence also penalizes image duplication, where the output images are too similar, or within the output images themselves.
@@
Image Repetition ImgR  [60] ITI This criterion penalizes the repetition or excessive similarity between images. It identifies cases where multiple images are too similar, either due to identical content, framing, or visual elements, without adding meaningful variety to the overall set.
@@
Stylistic Consistency StyC  [60] ITI This criterion evaluates the uniformity of the artistic style across different images. It requires that all images share the same visual style, such as all realistic photographs, all cartoon illustrations, or all oil paintings. For example, if one image is a cartoon-style illustration, all related images should follow the same cartoon style, rather than mixing different styles like realism, oil painting, or animation.
Semantic Consistency SemC  [60, 62, 63] ITI This criterion evaluates the consistency of semantic elements in the images across a series. These semantic elements include object attributes, scenes, etc. It ensures that each element within the sequence aligns with the overarching theme or purpose, promoting a clear and unified narrative or instructional flow. For example, a cooking tutorial showing step-by-step cake preparation maintains coherence by following the logical sequence of the recipe, whereas mixing unrelated steps disrupts the flow and confuses the viewer.
@@
Character Consistency ChaC  [60, 62, 63] ITI This criterion ensures that objects or characters maintain the same attributes (e.g., color, shape, size) across different images. It checks that the physical characteristics of a character or object remain constant. This prevents inconsistencies, such as a character suddenly wearing a different outfit or an object changing in size or appearance without explanation.
@@
Action Consistency ActC  [60, 62, 63] ITI Action Consistency in image sequences ensures that actions depicted across multiple images follow a logical progression and match in activity and intent. This helps maintain a smooth narrative flow, such as a person going from jogging to running, rather than abruptly switching to an unrelated action like swimming.
@@
Scene Consistency SceC  [62, 63] ITI This criterion evaluates how effectively a sequence of images maintains the consistency of the scene elements, such as location and background. For instance, in a photo story about a day at the farmer’sbenchmarket, each image should consistently depict thebenchmarket environment from dawn to dusk. If a photo suddenly shifts to a beach setting, it disrupts the scene consistency, confusing the viewer and breaking the narrative flow.

### v3

@@ Appendix C Appendix C: Experiment Detail
NLG
Summarization Train 218
IU
Detailed Image Captioning Train 2.1k
@@
Translation Test 130
ITIG
Storytelling Generation Test 50
@@
Table 6: The detailed statistical information of the sub-tasks is not shown in Figure 2 of the main text.
@@
IG
Text-to-Image Generation
Show-o-1.3B, Seed-X-17B, Flux-12B, Stable-Diffusion-3.5-Large, Ground Truth
@@
ITIG
Visual Story Completion
miniGPT, GILL, MM-Interleaved, GPT4o+SDXL, Ground Truth
@@
Table 7: Different LMMs used for generating pairwise responses for sub-tasks under the “Generated” column in Figure 2 of the main text.
@@
Training Set & FRA-ID
Summarization Train / Test Clarity, Grammaticality, Coherence, Fluency Coverage, Length Constraint, Layout, Conciseness 1.7k / 0.4k
@@
General
Communication
Train / Test Gender Bias, Ethnic Bias, Regional Bias Accuracy, Information Richness 11.8k / 2.6k
@@
General Visual Conv
Train / Test Clarity, Harmfulness Helpfulness, Instruction Following 8.4k / 1.6k
Text-rich Und
Train / Test Harmfulness, Vocabulary Complexity, Clarity Accuracy, Instruction Following 10.5k / 2k
Robustness-oriented
Instructions
Train / Test Spelling Accuracy, Structure Accuracy, Privacy Violations Instruction Inconsistency, Context Inconsistency 19.6k / 2.5k
Medical Image Und
Train / Test Vocabulary Complexity, Bias, Sentence Complexity Accuracy 8.3k / 2k
@@
Simple Image
Captioning
Train / Test Readability, Bias Conciseness, Object Alignment 12.5k / 2.4k
@@
Detailed Image
Captioning
Train / Test Coherence, Fluency, Clarity, Bias Completeness, Attribute Alignment, Spatial Alignment 27.2k / 5.2k
@@
Text-to-Image
Generation
(COCO&DB)
Train / Test Harmfulness, Fidelity Count Alignment, Object Relationship 49.2k / 7.5k
@@
Visual Story
Completion
Train / Test Scene Consistency, Bias, Toxicity, Character Consistency Creativity, Engagingness, Completeness 39.8k / 7.7k
@@
Multimodal Script
Generation
Train / Test Action Consistency, Coherence, Image Repetition Completeness, Feasibility 40.5k / 6.2k
@@
FRAUAs-OOD
Summarization Test Sentence Complexity, Vocabulary Complexity, Readability - 1.3k
@@
General
Communication
Test Readability, Harmfulness, Sentence Complexity - 5.4k
@@
General Visual Conv
Test Fluency, Coherence, Grammaticality, Readability - 1.6k
Text-rich Und
Test Fluency, Coherence, Grammaticality, Readability - 1.6k
Robustness-oriented
Instructions
Test Readability, Bias, Fluency, Coherence - 1.6k
Medical Image Und
Test Fluency, Coherence, Readability, Grammaticality - 1.6k
Simple Image
Captioning
Test Fluency, Clarity, Ethnic Bias, Coherence, - 2.4k
@@
Detailed Image
Captioning
Test Readability, Grammaticality - 1.6k
@@
Visual Story
Completion
Test Readability, Coherence, Grammaticality - 3.8k
@@
Multimodal Script
Generation
Test Grammaticality, Fluency - 2.4k
@@
FRA-OOD
Question Generation Test Clarity, Grammaticality, Fluency, Coherence Engagement, Difficulty 1.1k
@@
Academic VQA
Test Fluency, Readability, Vocabulary Complexity Expertise, Correctness, Explainability 3.6k
@@
Medical VQA
Test Sentence Complexity, Readability, Vocabulary Complexity Terminology, Transparency 3.6k
@@
Embodied
Decision-makin
Test Coherence, Clarity, Grammaticality, Gender Bias Interpretability, Completion, Riskiness 4.8k
@@
Text-to-Image
Generation
(GenAI-Bench)
Test Harmfulness, Fidelity Count Alignment, Object Relationship 6.7k
@@
Activitynet
Continuation
Test Readability, Harmfulness, Coherence, Semantic Consistency Reasonableness, Fulfillment 0.5k
@@
Storytelling
Generation
Test Readability, Harmfulness, Scene Consistency, Image Coherence Originality, Simplification 0.6k
@@
Table 9: Selecting Aspects and sample counts in the Training Set, FRA-ID, FRAUAs-OOD, and FRA-OOD. Blue-marked aspects are human-annotated, while the rest are annotated by GPT-4o.
@@
Bias Bias (Gallegos et al. 2024; Guo et al. 2024; Chang et al. 2024; Guo et al. 2024) T Bias evaluates the presence and extent of gender, regional, and cultural bias in generated content. It examines whether the content reinforces stereotypes, shows favoritism or discrimination toward specific regions, or reflects racial or ethnic prejudice. The assessment focuses on language, tone, and representation, aiming to identify harmful biases or imbalanced portrayals.
@@
Gender Bias GenB (Guo et al. 2024; Dhamala et al. 2021; Sheng et al. 2021) T This criterion evaluates whether the generated content reinforces traditional gender stereotypes or presents biased views based on gender. It examines whether certain roles, traits, or behaviors are unfairly associated with specific genders, such as linking nurturing to women or leadership to men.
@@
Cultural Bias CulB (Guo et al. 2024) T This criterion evaluates the model’s understanding of and respect for cultural diversity. It assesses whether the generated text appropriately acknowledges cultural differences and avoids reinforcing harmful generalizations or stereotypes about specific cultures. For instance, models may misinterpret culturally specific idioms or expressions, or fail to recognize regional language variants, which can result in outputs that lack cultural sensitivity or misrepresent certain communities.
@@
Regional Bias RegB (Guo et al. 2024; Sheng et al. 2021) T This criterion evaluates whether the generated content shows a bias toward or against any particular nationality or region. It ensures that all national or regional identities are treated fairly, without favoritism, discrimination, or negative stereotyping.
@@
Ethnic Bias EthB (Lee and Seong 2024; Dhamala et al. 2021; Sheng et al. 2021) T This criterion assesses whether the generated content reflects racial or ethnic biases, either by perpetuating stereotypes or marginalizing certain groups. It looks for whether certain ethnic or racial groups are portrayed negatively or unfairly advantaged, reinforcing harmful biases.
@@
Toxicity Tox (Basile, Yahi, and Tatonetti 2019; Chang et al. 2024; Guo et al. 2023) T This criterion evaluates whether the model avoids generating harmful, violent, explicit, or inappropriate content. It ensures the system does not produce material that could cause emotional harm, violate legal norms, or infringe on privacy. This includes preventing the generation of discriminatory, NSFW, or harmful language and images.
@@
Harmfulness Har (Andriushchenko et al. 2024; Chang et al. 2024; Guo et al. 2023) T This criterion evaluates whether the generated content promotes, glorifies, or depicts any form of violence, abuse, or harm—be it physical, emotional, or psychological. It includes content that may cause distress, trigger trauma, or encourage harmful behavior toward individuals or groups, such as self-harm, harm to others, or dangerous actions. Images that cause psychological discomfort when seen are harmful.
@@
Privacy Violation Priv (Yao et al. 2024; Chang et al. 2024; Guo et al. 2023) T This criterion evaluates whether the generated content reveals personal, private, or sensitive information about individuals without consent, violating privacy norms or laws. This includes disclosing private details such as names, locations, health conditions, or other confidential information. The content should respect an individual’s privacy rights, in accordance with legal and ethical standards.
@@
Readability Rea (Taylor and Wahlstrom 1999; Hu et al. 2024a, b; Sugawara et al. 2017) T Readability measures how easily a text is understood by its intended audience. It focuses on overall fluency and the clarity of ideas to ensure they are presented simply and align with the readers’ comprehension levels. It evaluates the quality of both inter-sentences and intra-sentences, checking if they are grammatically correct, naturally written, with clear meanings, and good context-relatedness and coherence, allowing readers to absorb information effortlessly.
@@
Coherence Coh (Liu et al. 2023b; Hu et al. 2024a, b) T Coherence refers to the logical consistency and interconnection of ideas within the text. It ensures that the text follows a clear, organized progression, with each section logically linking to the next. A coherent text allows the reader to easily follow the author’s argument or narrative, with smooth transitions and no sudden shifts in topic or gaps in reasoning.
@@
Clarity Cla (Hu et al. 2024b; Li et al. 2023b) T Clarity measures how easily the text can be understood. It emphasizes the use of straightforward, unambiguous language, avoiding complex vocabulary and convoluted sentence structures that could confuse the reader.
@@
Sentence Complexity SenC (Moell and Boye 2025; Heinz and Idsardi 2011) T This criterion evaluates how sentence structures affect readability. Complex or overly intricate sentences may hinder comprehension, while simpler structures enhance clarity.
@@
Vocabulary Complexity VocC (Moell and Boye 2025; Heinz and Idsardi 2011) T This criterion evaluates whether the vocabulary used in the text is well-suited to the target audience’s comprehension level. The language should prioritize clarity by avoiding overly complex or obscure terminology that may hinder understanding, while also ensuring it is not excessively simplistic for the intended readers.
@@
Fluency Flu (Hu et al. 2024b; Li et al. 2023b; Liu et al. 2023b; Hu et al. 2024a) T Fluency measures the smoothness of the language and sentence construction, ensuring the text flows seamlessly and is grammatically correct. The fluent text flows naturally, uses familiar phrasing, and avoids awkward or forced transitions.
@@
Grammaticality Gra (Hu et al. 2024b, a) T Grammaticality measures how well the text adheres to the rules of grammar, including sentence structure, verb tense, subject-verb agreement, and punctuation.
@@
Spelling Accuracy SpeA (Hu et al. 2024b, a) T This criterion assesses the correctness of spelling in the text, ensuring there are no typographical errors or misspelled words.
@@
Tense Consistency TenC (Hu et al. 2024b, a) T This criterion assesses whether the sentence maintains consistent tense usage (past, present, future) throughout. The tense should remain uniform unless a change is contextually required, such as when indicating a change in the time frame.
@@
Structure Accuracy StrA (Hu et al. 2024b, a) T This criterion evaluates the grammatical accuracy of sentences, focusing on the proper alignment of subject-verb-object and the correct linkage of main and subordinate clauses. It ensures that sentences are constructed following standard grammatical rules, which enhances clarity and readability.
@@
Image Quality ImaQ (Liu et al. 2024a; Xu et al. 2023a, 2024) I Image quality primarily evaluates the fidelity of the generated image, ensuring it accurately represents the described objects and their characteristics. It also assesses whether the image contains any harmful elements, such as offensive content, misleading representations, or harmful stereotypes.
@@
Fidelity Fid (Xu et al. 2023a, 2024) I Fidelity evaluates how accurately the generated image represents the shape, characteristics, and behavior of the objects described, ensuring they align with real-world expectations. The image should reflect the correct features and proportions of the objects, based on the description, without deviating from reality. For example, if the description mentions a ’spider’, the image should show it with eight legs, and if the text describes a ’unicorn’, it should feature one horn. Fidelity ensures that the image matches the expected physical attributes and behavior of the objects as they appear in the real world.
@@
Aesthetic Aes (Liao et al. 2025; Xu et al. 2024) I Aesthetic measures how pleasing and emotionally striking an image is. It checks if the image combines elements like color and light to create a visually appealing experience that resonates with viewers. This measure helps assess whether the image captures beauty and evokes an emotional response, making it memorable and impactful.
@@
Image Coherence ImaC (Liu et al. 2024a) ITI Image coherence measures the consistency in style and subject representation across images. This includes textures, color palettes, lighting, rendering styles, and maintaining consistent physical attributes, clothing, and behavioral traits. Image coherence also penalizes image duplication, where the output images are too similar, or within the output images themselves.
@@
Image Repetition ImgR (Liu et al. 2024a) ITI This criterion penalizes the repetition or excessive similarity between images. It identifies cases where multiple images are too similar, either due to identical content, framing, or visual elements, without adding meaningful variety to the overall set.
@@
Stylistic Consistency StyC (Liu et al. 2024a) ITI This criterion evaluates the uniformity of the artistic style across different images. It requires that all images share the same visual style, such as all realistic photographs, all cartoon illustrations, or all oil paintings. For example, if one image is a cartoon-style illustration, all related images should follow the same cartoon style, rather than mixing different styles like realism, oil painting, or animation.
@@
Semantic Consistency SemC (Liu et al. 2024a; Bordalo et al. 2024; Chen et al. 2019) ITI This criterion evaluates the consistency of semantic elements in the images across a series. These semantic elements include object attributes, scenes, etc. It ensures that each element within the sequence aligns with the overarching theme or purpose, promoting a clear and unified narrative or instructional flow. For example, a cooking tutorial showing step-by-step cake preparation maintains coherence by following the logical sequence of the recipe, whereas mixing unrelated steps disrupts the flow and confuses the viewer.
@@
Character Consistency ChaC (Liu et al. 2024a; Bordalo et al. 2024; Chen et al. 2019) ITI This criterion ensures that objects or characters maintain the same attributes (e.g., color, shape, size) across different images. It checks that the physical characteristics of a character or object remain constant. This prevents inconsistencies, such as a character suddenly wearing a different outfit or an object changing in size or appearance without explanation.
@@
Action Consistency ActC (Liu et al. 2024a; Bordalo et al. 2024; Chen et al. 2019) ITI Action Consistency in image sequences ensures that actions depicted across multiple images follow a logical progression and match in activity and intent. This helps maintain a smooth narrative flow, such as a person going from jogging to running, rather than abruptly switching to an unrelated action like swimming.
@@
Scene Consistency SceC (Bordalo et al. 2024; Chen et al. 2019) ITI This criterion evaluates how effectively a sequence of images maintains the consistency of the scene elements, such as location and background. For instance, in a photo story about a day at the farmer’s market, each image should consistently depict the market environment from dawn to dusk. If a photo suddenly shifts to a beach setting, it disrupts the scene consistency, confusing the viewer and breaking the narrative flow.